OpenStudy (anonymous):

If 40.27 mL of 0.100 M NaOH is required to neutralize 0.314 g of an unknown monoprotic acid, what is the molecular weight of this unknown acid?

OpenStudy (anonymous):

first you write the equation: NaOH + HX -> NaX + H2O now lets start with basics: n=m/M c=n/V so n(NaOH) = c(NaOH)*V(NaOH) n(NaOH) = 40,27 * 10^-3 dm3 * 0,1 mol dm^3 n(NaOH) = ??? mol n(HX) / n(NaOH) = 1/1 =1 so n(HX)=n(NaOH) M(HX) = m(HX) / n(HX) M(HX) = 0,314 g / ???? mol